OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (9,-5); 4x-3y=2 How do I go about doing this?

zepdrix (zepdrix):

Hmmmm

zepdrix (zepdrix):

First get your equation into that nice fancy y-intercept.. form.. or whatever it's called. y=mx+b Then we'll notice something interesting :o

OpenStudy (anonymous):

-3y=4x+2?

zepdrix (zepdrix):

Woops you need to subtract 4x from both sides to move it to the right :)

OpenStudy (anonymous):

4x-3y=2-4x right?

zepdrix (zepdrix):

 Ok take a look at that and lemme know if it makes sense.